Problem 5-2 Calculating Future Values [LO1]

For each of the following, compute the future value: (Do not round intermediate calculations and round your answers to 2 decimal places, e.g., 32.16.)

  

Present Value Years Interest Rate Future Value
$ 2,150 11      13 % $   
8,552 7      9   
74,355 14      12   
181,796 8      6   
Answer #1
FV = Future Value
PV = Present Value
r = rate of interest
n= no of period
a) FV = PV (1 + r )n
FV = 2150*(1+13%)^11
FV = 8247.10
b) FV = PV (1 + r )n
FV = 8552*(1+9%)^7
FV = 15633.39
c) FV = PV (1 + r )n
FV = 74355*(1+12%)^14
FV = 363381.23
d) FV = PV (1 + r )n
FV = 181796*(1+6%)^8
FV = 289755.20